OpenStudy (anonymous):

Differentiate and simplify. f(y)=(\frac{ 2-y }{ 1+2y })^{4}

OpenStudy (anonymous):

\[f(y)=\left( \frac{ 2-y }{ 1+2y } \right)^{4}\]

OpenStudy (diamondboy):

we would use chain rule and quotient rule

OpenStudy (anonymous):

Yes, and I got this: \[f'(y)=\frac{ 4(2-y)^3(1+2y)^4-(2-y)^44(1+2y)^3(2) }{ (1+2y)^8 }\] I'm not sure if this is correct...

OpenStudy (diamondboy):

 dis was wat I got for one half of d question